OpenStudy (anonymous):

Evaluate sin[2 arccos(-8/17 )].

OpenStudy (anonymous):

\[\cos^{-1}(\frac{-8}{7})\approx2.06\] Can you solve it now?

OpenStudy (anonymous):

Remember the mnemonic SOH-CAH-TOA \[\sin\theta=\frac{opposite}{hypotenuse}\] \[\cos\theta=\frac{adjacent}{hypotenuse}\]

OpenStudy (mertsj):

Let x = arccos(-8/17) sin(2x)=2sinxcosx

OpenStudy (mertsj):



OpenStudy (mertsj):

\[2sinxcosx=2(\frac{15}{17})(\frac{-8}{17})=\frac{-240}{289}=\sin2(\arccos\frac{-8}{17})\]